# Transfer learning for multifidelity simulation-based inference in cosmology

### v1

@@ 3.2 CAMELS Multifield Dataset: SB28 suite
We repeat the experiments of Section 3.1 on the SB28 suite, this time performing inference over a 5-dimensional posterior. The very broad range of nuisance parameters, as well as the extra three cosmological parameters $\{n_{\textrm{s}},h,\Omega_{\textrm{b}}\}$, lead to a more challenging inference problem. Prior work has only explored inferring $\Omega_{\textrm{m}}$ and $\sigma_{8}$ from this dataset, and has found that the larger set of cosmological and astrophysical parameters leads to much weaker constraints on $\sigma_{8}$ (Ni et al., 2023). Ni et al. (2023) also indicated that the Hubble constant $h$ and the baryonic fraction $\Omega_{\textrm{b}}$ have minor effects on the simulations, indicating that these may be challenging to constrain. Note that the SB28 suite contains roughly double the number of simulations as the LH suite, enabling a more accurate ML-based reference model.
@@
Fig. 7 displays the headline metrics comparing the two training approaches with a baseline “accurate” model that used the entire SB28 suite training set. Again, the transfer learning models significantly outperform models trained from scratch, and the MTPP score of the transfer learning experiments are only surpassed when training a high-fidelity-only model with $N=12,800$ IllustrisTNG maps. Panels b) and c) in Fig. 7 present a similar pattern as in Section 3.1: all transfer learning models are more constraining and better calibrated than training from scratch (until $N=12,800$). However, fine-tuned models display (minor) calibration issues until $N\geq 800$. The uptick in calibration error $\mathcal{C}$ for the transfer learning approach is very minor and largely within errors, so we do not attempt to interpret it. Interestingly, all fine-tuned models are better calibrated in the SB28 experiments than in the LH experiments (note the different $y$-axis scales).
@@
The performance difference between high-fidelity-only models and transfer learning models is even larger than in Section 3.1. Depending on the exact MTPP performance desired, Fig. 7 indicates that pre-training on $N$-body simulations allows for a factor of $10$ to $15$ reduction in high-fidelity simulations to train an informative, well-calibrated model of the posterior.
@@
We found that none of the models could constrain $\Omega_{\textrm{b}}$ and $h$ far beyond the uniform prior. Given that this was a feature of the “accurate” baseline model, trained on $27720$ maps, we may conclude that this a genuine feature of the simulations, at least up to the resolving power of the CNN-NDE architecture used to perform inference. This is reflected by the FoM results in Fig. 7 c), which are significantly lower for all models than the FoM in the LH suite experiment. We present two examples of inference in Fig. 8, showing only $\{\Omega_{\textrm{m}},\sigma_{8},n_{\textrm{s}}\}$. This implicitly marginalises over the poorly constrained $\Omega_{\textrm{b}}$ and $h$. Again, we compare training from scratch with $N=800$ and fine-tuning with just $N=800$ maps against an “accurate” baseline. We present examples of inference of the full 5-dimensional posterior in Appendix C.
@@
Again, multifidelity transfer learning produces a model that significantly outperforms high-fidelity-only training. The difference in posterior quality is even more stark than in Fig. 5; the high-fidelity-only posteriors are very uninformative and fail to extract much useful cosmological information from the density maps. The overconfidence identified in Fig. 8 is apparent as a bias in the left panel of Fig. 5. On the other hand, the transfer learning approach recovers the key features of the accurate baseline posteriors, including both the location and width of the posterior contours. Both the “accurate” and transfer learning models yield a degeneracy between the amplitude of the matter density power spectrum $\sigma_{\textrm{8}}$ and the scalar spectral index $n_{\textrm{s}}$. This degeneracy is expected for two-point statistics at the very short scales probed in the CMD simulations, given both parameters have similar, difficult to distinguish marginal effects on the power spectrum. The results in Fig. 8 indicate that the non-linear effects probed by the CNN are insufficient to fully break this degeneracy.
@@
The larger performance gains reported here relative to the LH suite in Section 3.1 likely result from the more complex task of modelling a 5-dimensional posterior and marginalising over a larger set of astrophysical parameters. We tentatively conclude that transfer learning may perform even better in more challenging inference problems, particularly those involving higher-dimensional posteriors and a broader set of nuisance parameters. As the complexity of the target task increases, the value of incorporating prior knowledge through pre-training is likely to grow.
@@ 4 Discussion and Conclusions
In this study, we have demonstrated that leveraging multifidelity simulations can significantly reduce the number of expensive simulations required to perform cosmological inference with SBI. By pre-training a neural inference model on a large set of lower-fidelity dark matter only simulations, we were able to perform informative and well-calibrated inference on IllustrisTNG hydrodynamical simulations with $<1000$ high-fidelity dark matter maps. This is a substantial improvement over previous work, which had demonstrated that training neural compression algorithms with small datasets led to suboptimal compression and inference (Hermans et al., 2022; Park et al., 2025; Bairagi et al., 2025; Jeffrey et al., 2025). The relative simplicity of our framework makes this method broadly applicable across cosmology.
@@ Acknowledgements
AAS was supported by the STFC UCL Centre for Doctoral Training in Data Intensive Science (grant ST/W00674X/1) and by departmental and industry contributions. AAS was also supported by the A. G. Leventis Foundation educational grant scheme. DP was supported by a Swiss National Science Foundation (SNSF) grant, and by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”. BJ acknowledges support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1.
@@ Appendix A Further posterior comparisons
We present a further comparison between high-fidelity-only training, with $N=3200$, and the transfer learning approach for the LH suite 2-parameter inference problem in Fig. 9. Despite the factor of ×$\times 4$ increase in the number of high-fidelity maps used during training, the multifidelity approach yields significantly tighter posteriors that better match the “accurate” model. In addition, the right-most panel in Fig. 9 is suggestive of the overconfidence issue identified in Section 3.1; high-fidelity-only models trained with fewer than $N=6400$ IllustrisTNG maps exhibit a large degree of overconfidence, per Fig. 4.
@@ Appendix B Probing model performance
Figure 10: Comparing inference results on the IllustrisTNG LH suite for various experiments. We reproduce the results from Fig. 3 for a) MTPP (note that the y-axis scale has been shortened for enhanced visualisation) and b) FoM. We also show results from experiments: pre-training and fine-tuning without any paired data (orange); training a neural compression model with $N$ IllustrisTNG maps and then freezing the CNN compression to train an NDE with the full ($N=12000$) LH training suite (red); and the performance of models only pre-trained on $N$-body simulations (black, corresponding to $N=0$ IllustrisTNG maps).
@@
Figure 11: Comparing transfer learning results on the IllustrisTNG LH suite using $M_{\textrm{cdm}}$ (dot-dashed lines) and $M_{\textrm{tot}}$ (solid lines). We reproduce the $M_{\textrm{cdm}}$ results from Fig. 3 for a) MTPP and b) calibration error. Pre-training on $N$-body simulations gives near equivalent improvements over training with only high-fidelity maps when performing inference on $M_{\textrm{tot}}$ fields.
@@
We performed a range of experiments to better understand the model performance. The results of these experiments are summarised in Fig. 10. We explored the impact of the paired aspect of the multifidelity simulation suite (i.e. that each simulation in the lower-fidelity $N$-body suite is paired with a high-fidelity simulation with identical cosmological parameters and initial conditions). This could potentially improve performace, perhaps due to implicit memorisation of the (pre-)training data (a well-studied phenomenon in deep learning, see e.g. Yeom et al., 2018; Carlini et al., 2019, 2023). We tested this by ensuring different (i.e. unpaired) cosmologies where used during pre-training and fine-tuning, and found that pairing had no discernable impact on performance.
@@
We further investigated the neural network inference model by splitting it into two components: the neural compression performed by the CNN, and the density estimation of the NDE. In order to better disambiguate the role of each component, we took the best transfer learning models from Section 3.1 and froze the CNN. This fixed the summary statistics that were extracted from the dark matter density maps for a given transfer learning size $N$. We then retrained the NDE with the entire training dataset ($N=12000$) with the frozen neural compression model. The resulting performance is shown in red in Fig. 10. We found that for very low compression training dataset size ($N=200$), retraining the NDE with much more data led to improved MTPP performance even without more informative summary statistics. This presents more evidence that for very low fine-tuning dataset sizes, the NDE is an important bottleneck of performance. However, as the fine-tuning dataset size increases, the model performance reverts to the standard transfer learning baseline. This indicates that for larger fine-tuning datasets, the difference between an “accurate” model trained on the full LH training suite and the transfer learning models is driven by slightly poorer neural compression. These results suggest that while pre-training on $N$-body simulations encourages highly informative summaries, there are likely subtle differences in the high-fidelity IllustrisTNG simulations (that are useful for slightly improving cosmological constraints) which the CNN fails to discover during fine-tuning.
@@
Fig. 10 also shows the $N=0$ transfer learning case, where only the $N$-body simulation pre-training is performed and no high-fidelity maps are used. The extremely poor performance indicates that there are significant differences between the different simulation fidelities, and a fine-tuning step is necessary.
@@
Finally, we test whether transfer learning for (the more observationally important) total matter density $M_{\textrm{tot}}$ behaves any differently. A key concern is that the improved performance from transfer learning could be largely due to the strong similarity between the IllustrisTNG dark matter density $M_{\textrm{cdm}}$ and that of dark matter-only $N$-body simulations. Figure 11 demonstrates that multifidelity transfer learning performs just as well when fine-tuning on $M_{\textrm{tot}}$.
@@
We compare the results from Section 3.1 with inference on the $M_{\textrm{tot}}$ field with an identical methodology. We find that transfer learning still leads to up to an order-of-magnitude reduction in the number of high-fidelity maps required to train an accurate, trustworthy inference model, compared with high-fidelity-only training.
@@
The small downward shift of all $M_{\textrm{tot}}$ inference performance curves (“accurate”, transfer learning and high-fidelity-only) on the MTPP metric from Section 3.1a indicates that inference using the $M_{\textrm{tot}}$ maps is slightly more challenging. However, there is also a slightly larger gap between the “accurate” $M_{\textrm{tot}}$ model and transfer learned models (and low $N$ high-fidelity-only models) compared with inference results on $M_{\textrm{tot}}$. This suggests that: i) some features in $M_{\textrm{tot}}$ require a large number of training maps ($N>6400$) for the CNN to learn to extract, more-so than in the $M_{\textrm{cdm}}$ case, and ii) $N$-body pre-training gives slightly less informative features than in the $M_{\textrm{cdm}}$ case, perhaps for similar reasons as i).
@@ Appendix C Unconstrained parameters in SB28
The posterior estimation models in Section 3.2 were trained to perform 5-dimensional inference on $\{\Omega_{\textrm{m}},\sigma_{8},n_{\textrm{s}},h,\Omega_{\textrm{b}}\}$. However, we found that $h$ and $\Omega_{\textrm{b}}$ could not be constrained by the data (or, partially, by the CNN-NDE architecture). Here we present some more details on these unconstrained parameters.
Figure 12 shows the posterior sample ensemble mean MSE for two cosmological parameters: $\sigma_{8}$ and $\Omega_{\textrm{b}}$. The posterior recovery of $\sigma_{8}$ behaves similarly to Section 3.1, with very good performance relative to the “accurate” baseline. On the other hand, we find that pre-training on $N$-body simulations leads to no improvement over the high-fidelity-only training approach for $\Omega_{\textrm{b}}$. In one sense this is expected: $N$-body simulations do not provide a strong probe of how $\Omega_{\textrm{b}}$ affects dark matter maps (beyond the initial matter power spectrum), and so there should not be much direct transfer of knowledge. In addition, Fig. 13 demonstrates that there is little constraining information on $\Omega_{\textrm{b}}$ in the dark matter density maps.
@@
Two examples of 5-dimensional posterior inference are given in Fig. 13. In the first example, none of the models can constrain $h$ and $\Omega_{\textrm{b}}$ much beyond the uniform prior. We can highlight two key qualitative features: the fine-tuned approach gives significantly better agreement with the baseline than training from scratch, and it is statistically consistent with the accurate posterior. In the second example, the “accurate” posterior gives a (weak) constraint on $\Omega_{\textrm{b}}$, while the other models fail to provide any constraints. A small but not insignificant fraction of the inferred posteriors follows this second pattern, which is consistent with the minor improvement in constraining power of $\Omega_{\textrm{b}}$ shown in Fig. 12. We found that these examples tended to coincide with extreme cosmologies (at the boundaries of the prior volume), and particularly for large values of $\Omega_{\textrm{b}}$ as is the case in Fig. 13. We reserve a more systematic analysis as a potential avenue for future work.

### v2

@@ 3.2 CAMELS Multifield Dataset: SB28 suite
We repeat the experiments of Section˜3.1 on the SB28 suite, this time performing inference over a 5-dimensional posterior. The very broad range of nuisance parameters, as well as the extra three cosmological parameters $\{n_{\textrm{s}},h,\Omega_{\textrm{b}}\}$, lead to a more challenging inference problem. Prior work has only explored inferring $\Omega_{\textrm{m}}$ and $\sigma_{8}$ from this dataset, and has found that the larger set of cosmological and astrophysical parameters leads to much weaker constraints on $\sigma_{8}$ (Ni et al., 2023). Ni et al. (2023) also indicated that the Hubble constant $h$ and the baryonic fraction $\Omega_{\textrm{b}}$ have minor effects on the simulations, indicating that these may be challenging to constrain. Note that the SB28 suite contains roughly double the number of simulations as the LH suite, enabling a more accurate ML-based reference model.
@@
Fig.˜7 displays the headline metrics comparing the two training approaches with a baseline “accurate” model that used the entire SB28 suite training set. Again, the transfer learning models significantly outperform models trained from scratch, and the MTPP score of the transfer learning experiments are only surpassed when training a high-fidelity-only model with $N=12,800$ IllustrisTNG maps. Panels b) and c) in Fig.˜7 present a similar pattern as in Section˜3.1: all transfer learning models are more constraining and better calibrated than training from scratch (until $N=12,800$). However, fine-tuned models display (minor) calibration issues until $N\geq 800$. The uptick in calibration error $\mathcal{C}$ for the transfer learning approach is very minor and largely within errors, so we do not attempt to interpret it. Interestingly, all fine-tuned models are better calibrated in the SB28 experiments than in the LH experiments (note the different $y$-axis scales).
@@
The performance difference between high-fidelity-only models and transfer learning models is even larger than in Section˜3.1. Depending on the exact MTPP performance desired, Fig.˜7 indicates that pre-training on $N$-body simulations allows for a factor of $10$ to $15$ reduction in high-fidelity simulations to train an informative, well-calibrated model of the posterior.
@@
We found that none of the models could constrain $\Omega_{\textrm{b}}$ and $h$ far beyond the uniform prior. Given that this was a feature of the “accurate” baseline model, trained on $27720$ maps, we may conclude that this a genuine feature of the simulations, at least up to the resolving power of the CNN-NDE architecture used to perform inference. This is reflected by the FoM results in Fig.˜7 c), which are significantly lower for all models than the FoM in the LH suite experiment. We present two examples of inference in Fig.˜8, showing only $\{\Omega_{\textrm{m}},\sigma_{8},n_{\textrm{s}}\}$. This implicitly marginalises over the poorly constrained $\Omega_{\textrm{b}}$ and $h$. Again, we compare training from scratch with $N=800$ and fine-tuning with just $N=800$ maps against an “accurate” baseline. We present examples of inference of the full 5-dimensional posterior in Appendix˜C.
@@
Again, multifidelity transfer learning produces a model that significantly outperforms high-fidelity-only training. The difference in posterior quality is even more stark than in Fig.˜5; the high-fidelity-only posteriors are very uninformative and fail to extract much useful cosmological information from the density maps. The overconfidence identified in Fig.˜8 is apparent as a bias in the left panel of Fig.˜5. On the other hand, the transfer learning approach recovers the key features of the accurate baseline posteriors, including both the location and width of the posterior contours. Both the “accurate” and transfer learning models yield a degeneracy between the amplitude of the matter density power spectrum $\sigma_{\textrm{8}}$ and the scalar spectral index $n_{\textrm{s}}$. This degeneracy is expected for two-point statistics at the very short scales probed in the CMD simulations, given both parameters have similar, difficult to distinguish marginal effects on the power spectrum. The results in Fig.˜8 indicate that the non-linear effects probed by the CNN are insufficient to fully break this degeneracy.
@@
The larger performance gains reported here relative to the LH suite in Section˜3.1 likely result from the more complex task of modelling a 5-dimensional posterior and marginalising over a larger set of astrophysical parameters. We tentatively conclude that transfer learning may perform even better in more challenging inference problems, particularly those involving higher-dimensional posteriors and a broader set of nuisance parameters. As the complexity of the target task increases, the value of incorporating prior knowledge through pre-training is likely to grow.
@@ 4 Discussion and Conclusions
In this study, we have demonstrated that leveraging multifidelity simulations can significantly reduce the number of expensive simulations required to perform cosmological inference with SBI. By pre-training a neural inference model on a large set of lower-fidelity dark matter only simulations, we were able to perform informative and well-calibrated inference on IllustrisTNG hydrodynamical simulations with <1000<1000 high-fidelity dark matter maps. This is a substantial improvement over previous work, which had demonstrated that training neural compression algorithms with small datasets led to suboptimal compression and inference (Hermans et al., 2022; Park et al., 2025; Bairagi et al., 2025; Jeffrey et al., 2025). The relative simplicity of our framework makes this method broadly applicable across cosmology.
@@ Acknowledgements
AAS was supported by the STFC UCL Centre for Doctoral Training in Data Intensive Science (grant ST/W00674X/1) and by departmental and industry contributions. AAS was also supported by the A. G. Leventis Foundation educational grant scheme. DP was supported by a Swiss National Science Foundation (SNSF) grant, and by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”. AMGF is grateful to support from the UPFLOW project, funded by the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No 101001601). NJ and BJ acknowledge support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1.
@@ Appendix A Further posterior comparisons
We present a further comparison between high-fidelity-only training, with $N=3200$, and the transfer learning approach for the LH suite 2-parameter inference problem in Fig.˜9. Despite the factor of $\times 4$ increase in the number of high-fidelity maps used during training, the multifidelity approach yields significantly tighter posteriors that better match the “accurate” model. In addition, the right-most panel in Fig.˜9 is suggestive of the overconfidence issue identified in Section˜3.1; high-fidelity-only models trained with fewer than $N=6400$ IllustrisTNG maps exhibit a large degree of overconfidence, per Fig.˜4.
@@ Appendix B Probing model performance
Figure 10: Comparing inference results on the IllustrisTNG LH suite for various experiments. We reproduce the results from Fig.˜3 for a) MTPP (note that the y-axis scale has been shortened for enhanced visualisation) and b) FoM. We also show results from experiments: pre-training and fine-tuning without any paired data (orange); training a neural compression model with $N$ IllustrisTNG maps and then freezing the CNN compression to train an NDE with the full ($N=12000$) LH training suite (red); and the performance of models only pre-trained on $N$-body simulations (black, corresponding to $N=0$ IllustrisTNG maps).
@@
Figure 11: Comparing transfer learning results on the IllustrisTNG LH suite using $M_{\textrm{cdm}}$ (dot-dashed lines) and $M_{\textrm{tot}}$ (solid lines). We reproduce the $M_{\textrm{cdm}}$ results from Fig.˜3 for a) MTPP and b) calibration error. Pre-training on $N$-body simulations gives near equivalent improvements over training with only high-fidelity maps when performing inference on $M_{\textrm{tot}}$ fields.
Figure 12: The convergence behaviour of our approaches as a function of dataset size for the LH suite. The legend shows which learning rate scheduler (cyclic or exp) was used, which we found had a minor effect on asymptotic behaviour. We reproduce the $M_{\textrm{cdm}}$ results from Fig.˜3 and extend the experiments to run all the way to $N=12000$ (blue and green lines). We also report the results of transfer learning using the high learning rate (LR) and cyclic scheduler settings used for the “Accurate” baseline model (orange line).
@@
We performed a range of experiments to better understand the model performance. The results of these experiments are summarised in Fig.˜10. We explored the impact of the paired aspect of the multifidelity simulation suite (i.e. that each simulation in the lower-fidelity $N$-body suite is paired with a high-fidelity simulation with identical cosmological parameters and initial conditions). This could potentially improve performace, perhaps due to implicit memorisation of the (pre-)training data (a well-studied phenomenon in deep learning, see e.g. Yeom et al., 2018; Carlini et al., 2019, 2023). We tested this by ensuring different (i.e. unpaired) cosmologies where used during pre-training and fine-tuning, and found that pairing had no discernable impact on performance.
@@
We further investigated the neural network inference model by splitting it into two components: the neural compression performed by the CNN, and the density estimation of the NDE. In order to better disambiguate the role of each component, we took the best transfer learning models from Section˜3.1 and froze the CNN. This fixed the summary statistics that were extracted from the dark matter density maps for a given transfer learning size $N$. We then retrained the NDE using the high-fidelity-only approach on the entire training dataset ($N=12000$) with the frozen neural compression model. The resulting performance is shown in red in Fig.˜10. The frozen CNN results (red) are produced by taking the CNN from the transfer-learning approach (green), freezing its weights, and retraining the NDE with the full high-fidelity dataset. We observe that at low $N$ (e.g., $N=200$), the model with a frozen CNN but fully retrained NDE (red) outperforms the standard transfer learning baseline (green), suggesting that the NDE limits transfer learning performance in the low $N$ regime. As $N$ increases the baseline transfer learning model catches up, and by $N=2400$ both approaches perform similarly, indicating that the NDE is no longer a performance bottleneck. This convergence implies that the main limitation of transfer learning at higher $N$ (i.e., $N>2400$) is due to slightly suboptimal CNN-based compression. These results suggest that while pre-training on $N$-body simulations encourages highly informative summaries, there are likely subtle differences in the high-fidelity IllustrisTNG simulations (that are useful for slightly improving cosmological constraints) which the CNN fails to discover during fine-tuning.
@@
Fig.˜10 also shows the $N=0$ transfer learning case, where only the $N$-body simulation pre-training is performed and no high-fidelity maps are used. The extremely poor performance indicates that there are significant differences between the different simulation fidelities, and a fine-tuning step is necessary.
@@
Next, we test whether transfer learning for (the more observationally important) total matter density $M_{\textrm{tot}}$ behaves any differently. A key concern is that the improved performance from transfer learning could be largely due to the strong similarity between the IllustrisTNG dark matter density $M_{\textrm{cdm}}$ and that of dark matter-only $N$-body simulations. Figure˜11 demonstrates that multifidelity transfer learning performs just as well when fine-tuning on $M_{\textrm{tot}}$.
@@
We compare the results from Section˜3.1 with inference on the $M_{\textrm{tot}}$ field with an identical methodology. We find that transfer learning still leads to up to an order-of-magnitude reduction in the number of high-fidelity maps required to train an accurate, trustworthy inference model, compared with high-fidelity-only training.
@@
The small downward shift of all $M_{\textrm{tot}}$ inference performance curves (“accurate”, transfer learning and high-fidelity-only) on the MTPP metric from Section˜3.1a indicates that inference using the $M_{\textrm{tot}}$ maps is slightly more challenging. However, there is also a slightly larger gap between the “accurate” $M_{\textrm{tot}}$ model and transfer learned models (and low $N$ high-fidelity-only models) compared with inference results on $M_{\textrm{tot}}$. This suggests that: i) some features in $M_{\textrm{tot}}$ require a large number of training maps ($N>6400$) for the CNN to learn to extract, more-so than in the $M_{\textrm{cdm}}$ case, and ii) $N$-body pre-training gives slightly less informative features than in the $M_{\textrm{cdm}}$ case, perhaps for similar reasons as i).
Finally, we present results exploring the convergence properties of transfer learning and high-fidelity-only training in Fig.˜12. We extend the dataset sizes up to $N=12000$ for the high-fidelity-only and transfer learning approaches. The additional line (orange) shows the results from transfer learning when using the same learning rate and scheduler as the “accurate” reference models. We find that all models asymptote slightly below the “accurate” model baseline. In the high-fidelity-only approach, this is due to a difference in scheduler; as discussed in the text, we find a cyclic learning rate scheduler (“accurate”, red) slightly outperforms an exponential scheduler for large dataset sizes ($N>10000$). In the transfer learning case, using the same learning rate and scheduler as the “accurate” models (orange) gets closer to recovering the baseline at $N=12000$, but this leads to higher variance training with worse calibration at smaller dataset sizes. We therefore conclude that pre-training can lead to a very minor reduction in performance when the pre-training dataset is no larger than the target dataset, but that in all other instances it is desirable.
@@ Appendix C Unconstrained parameters in SB28
The posterior estimation models in Section˜3.2 were trained to perform 5-dimensional inference on $\{\Omega_{\textrm{m}},\sigma_{8},n_{\textrm{s}},h,\Omega_{\textrm{b}}\}$. However, we found that $h$ and $\Omega_{\textrm{b}}$ could not be constrained by the data (or, partially, by the CNN-NDE architecture). Here we present some more details on these unconstrained parameters.
Figure˜13 shows the posterior sample ensemble mean MSE for two cosmological parameters: $\sigma_{8}$ and $\Omega_{\textrm{b}}$. The posterior recovery of $\sigma_{8}$ behaves similarly to Section˜3.1, with very good performance relative to the “accurate” baseline. On the other hand, we find that pre-training on $N$-body simulations leads to no improvement over the high-fidelity-only training approach for $\Omega_{\textrm{b}}$. In one sense this is expected: $N$-body simulations do not provide a strong probe of how $\Omega_{\textrm{b}}$ affects dark matter maps (beyond the initial matter power spectrum), and so there should not be much direct transfer of knowledge. In addition, Fig.˜14 demonstrates that there is little constraining information on $\Omega_{\textrm{b}}$ in the dark matter density maps.
@@
Two examples of 5-dimensional posterior inference are given in Fig.˜14. In the first example, none of the models can constrain $h$ and $\Omega_{\textrm{b}}$ much beyond the uniform prior. We can highlight two key qualitative features: the fine-tuned approach gives significantly better agreement with the baseline than training from scratch, and it is statistically consistent with the accurate posterior. In the second example, the “accurate” posterior gives a (weak) constraint on $\Omega_{\textrm{b}}$, while the other models fail to provide any constraints. A small but not insignificant fraction of the inferred posteriors follows this second pattern, which is consistent with the minor improvement in constraining power of $\Omega_{\textrm{b}}$ shown in Fig.˜13. We found that these examples tended to coincide with extreme cosmologies (at the boundaries of the prior volume), and particularly for large values of $\Omega_{\textrm{b}}$ as is the case in Fig.˜14. We reserve a more systematic analysis as a potential avenue for future work.